wifeofsosa:

when we solve an inequality and divide by a negative what do you do with the inequality sign?

jhonyy9:

-3x > 6 divide both sides by -3 so what will get ? @XioGonz

wifeofsosa:

a negative?

XioGonz:

\[-3x > 6\] \[\frac{ -3x }{ 3 }>\frac{ 6 }{ 3 }\] \[x<-2\]

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \[-3x > 6\] \[\frac{ -3x }{ 3 }>\frac{ 6 }{ 3 }\] \[x<-2\] \(\color{#0cbb34}{\text{End of Quote}}\) divide by minus 3 not plus 3

XioGonz:

See how the sign was flipped?

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \[-3x > 6\] \[\frac{ -3x }{ -3 }>\frac{ 6 }{ -3 }\] \[x<-2\] \(\color{#0cbb34}{\text{End of Quote}}\) divide by minus 3 not plus 3 \(\color{#0cbb34}{\text{End of Quote}}\) My bad Let me correct

XioGonz:

There

YSNJayy:

yeah i was about to say that the sign would flip

wifeofsosa:

i still dont understand

jimthompson5910:

Using @jhonyy9 's example, we can solve for x without dividing both sides by a negative number. To do this, we need to add 3x to both sides so that the -3 becomes positive. -3x > 6 -3x+3x > 6+3x 0x > 3x+6 0 > 3x+6 Then next, we subtract 6 from both sides. 0 > 3x+6 0-6 > 3x+6-6 -6 > 3x To isolate x fully, you need to divide both sides by 3 to undo the multiplication. -6 > 3x -6/3 > 3x/3 -2 > x So -2 is larger than x, which is the same as saying x is smaller than -2 In other words, -2 > x becomes x < -2 Side note: Dividing both sides by a positive number won't flip the inequality sign.

wifeofsosa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 Using @jhonyy9 's example, we can solve for x without dividing both sides by a negative number. To do this, we need to add 3x to both sides so that the -3 becomes positive. -3x > 6 -3x+3x > 6+3x 0x > 3x+6 0 > 3x+6 Then next, we subtract 6 from both sides. 0 > 3x+6 0-6 > 3x+6-6 -6 > 3x To isolate x fully, you need to divide both sides by 3 to undo the multiplication. -6 > 3x -6/3 > 3x/3 -2 > x So -2 is larger than x, which is the same as saying x is smaller than -2 In other words, -2 > x becomes x < -2 Side note: Dividing both sides by a positive number won't flip the inequality sign. \(\color{#0cbb34}{\text{End of Quote}}\) oh ok thank you

jimthompson5910:

No problem